OpenStudy (anonymous):

what is the product in simplest form? state any restrictions. x^2+7x+10/x+3 times x^2-3x-18/x^2+x-2

OpenStudy (gorv):

\[x \neq -3\]

OpenStudy (anonymous):

?does that mean the answer is -3?

OpenStudy (gorv):

\[x^2+x-2 \neq 0\]

OpenStudy (anonymous):

what?

OpenStudy (gorv):

no no it is just restriction

OpenStudy (gorv):

denominator should not be zero otherwise we will get infinity as ans

OpenStudy (anonymous):

im so confused you have to understand i have no clue what im doing.....im a bit of a dunce when it comes to mathematics.

OpenStudy (gorv):

jst a min

OpenStudy (gorv):

x^2+x-2=0 x^2+2x-x-2=0 x(x+2)-1(x+2)=0 (x-1)(x+2)=0

OpenStudy (gorv):

\[x \neq 1\] \[x \neq -2\]

OpenStudy (gorv):

our ans should not be infinity

OpenStudy (gorv):

if denominator equals to zero than numerator/zero=infinity so we need to avoid those values which makes denominator zero

OpenStudy (gorv):

@kitty_love

OpenStudy (anonymous):

sooo -1?

OpenStudy (gorv):

at x=1 x=-2 and x=-3 denominator equals to zero so avoid these three value

OpenStudy (anonymous):

ok so the restrictions are 1,-2,and 3?

OpenStudy (gorv):

x^2+7x+10= (x+5)(x+2) on factoring x^2-3x-18=(x-6)(x+3) on factoring

OpenStudy (gorv):

yeah and that is -3 not 3

OpenStudy (anonymous):

? so which is it?

OpenStudy (gorv):

wait i m solving

OpenStudy (gorv):

\[\frac{ (x+5)*(x-6) }{ x-2 }\]

OpenStudy (gorv):

this is ur final ans

OpenStudy (anonymous):

thank you so much. and thanks for being pationant.